### v1

@@ Abstract
High-definition 3D city maps underpin smart transportation, digital twins, and autonomous driving, where object level change detection across bi temporal LiDAR enables HD map maintenance, construction monitoring, and reliable localization. Classical DSM differencing and image based methods are sensitive to small vertical bias, ground slope, and viewpoint mismatch and yield cellwise outputs without object identity. Point based neural models and voxel encodings demand large memory, assume near perfect pre alignment, degrade thin structures, and seldom enforce class consistent association, which leaves split or merge cases unresolved and ignores uncertainty. We propose an object centric, uncertainty aware pipeline for city scale LiDAR that aligns epochs with multi resolution NDT followed by point to plane ICP, normalizes height, and derives a per location level of detection from registration covariance and surface roughness to calibrate decisions and suppress spurious changes. Geometry only proxies seed cross epoch associations that are refined by semantic and instance segmentation and a class constrained bipartite assignment with augmented dummies to handle splits and merges while preserving per class counts. Tiled processing bounds memory without eroding narrow ground changes, and instance level decisions combine 3D overlap, normal direction displacement, and height and volume differences with a histogram distance, all gated by the local level of detection to remain stable under partial overlap and sampling variation. On 15 representative Subiaco blocks the method attains 95.2% accuracy, 90.4% mF1, and 82.6% mIoU, exceeding Triplet KPConv by 0.2 percentage points in accuracy, 0.2 in mF1, and 0.8 in mIoU, with the largest gain on Decreased where IoU reaches 74.8% and improves by 7.6 points. The Subiaco 2025 HD LiDAR dataset and annotations are released at IEEE DataPort, and the implementation is available at GitHub.
Index Terms:
3D change detection, LiDAR, point clouds, urban mapping, uncertainty-aware registration
[…]
=-15pt
@@ I Introduction
Western Australia is advancing smart transportation, digital-twin, and autonomous driving deployments that depend on high-definition 3D city maps maintained at operational cadence [1, 2]. Within this context, change detection denotes the object-level identification of additions, removals, and volumetric increases or decreases across bi-temporal LiDAR maps, enabling HD-map maintenance, construction and asset monitoring, and timely updates for localization and planning in autonomy stacks [3, 4, 5, 6, 7]. LiDAR sensing further supports this task by providing dense geometry that separates long-term structural changes from transient dynamics such as vehicles and pedestrians [8, 9, 10]. Subiaco lies immediately west of Perth’s CBD and functions as a key transport hub with two arterial roads linking the areas and substantial daily commuting, which makes it a representative and consequential testbed [11, 12]. Addressing city-scale change detection in Subiaco is therefore directly aligned with Western Australia’s smart-city and autonomy objectives and motivates the problem tackled in this work [13].
Existing approaches to urban 3D change detection fall into height or digital surface model differencing, 2D image based change detection, and point based deep models [14, 15]. Height differencing requires sub decimeter co registration and uniform sampling [16, 17]. In dense street scenes small vertical biases and ground slope introduce systematic errors, curb lines and façade parallax generate false positives, and the output remains cell wise without object identity [18, 16]. Image based change detection cannot resolve vertical geometry and suffers from viewpoint mismatch with ground LiDAR [19, 20]. Point wise Siamese or triplet networks such as KPConv variants operate on cropped tiles with millions of points [21, 22, 23]. Runtime and memory grow with area, training assumes balanced labels, and predictions are per point rather than per object, which weakens Added and Removed reasoning when overlap is partial or occlusion changes sampling [24, 25]. Voxel or bird’s eye encodings reduce memory but erode thin structures and narrow ground changes, which forces a trade off between coverage and fidelity [26, 27]. Public benchmarks for 3D change detection cover limited extents or rely on aerial campaigns, and object level annotations across epochs are rare [28, 29]. For Subiaco, publicly accessible LiDAR captures are single epoch and not paired at consistent specification, and no object level multi epoch ground truth exists [28]. Many pipelines also assume perfect pre alignment and do not propagate registration uncertainty into decision thresholds, which leads to spurious changes near loop junctions or under canopy [30, 31, 32, 33, 34]. Methods seldom enforce class consistent association between epochs, so split and merge cases remain unresolved and per class counts drift at city scale [35, 36, 29]. These constraints limit transfer to operational HD map maintenance over a dense transport hub and expose the need for a city scale, object centric and uncertainty aware formulation.
@@
Figure 1: Overview of the Subiaco city-scale 3D change-detection pipeline. Top: acquisition/mapping (soft GNSS sync, ground segmentation, sequence construction, NDT $\rightarrow$ ICP). Bottom: analysis (cross-epoch association, semantic/object segmentation, and uncertainty-gated instance labels).
Our changing detection method is object centric and uncertainty aware, and it operates at city scale on bi-temporal LiDAR maps. An overview of the acquisition, mapping, and change-analysis stages is shown in Fig. 1. It begins with global alignment of the two epochs using multi-resolution NDT followed by point-to-plane ICP, after which both maps are height normalized and a per-location level-of-detection is derived from registration covariance and surface roughness. This step replaces the usual perfect pre-alignment assumption and supplies calibrated thresholds that suppress spurious changes along loop junctions and under canopy. We then generate geometry-only proxies to seed cross-epoch associations and perform semantic and instance segmentation for ground, building, vegetation, and mobile classes on single-epoch geometry. Class-constrained bipartite matching refines correspondences while handling split and merge cases through an augmented assignment with dummies, which prevents drift in per-class counts and enforces label consistency across time. The city is partitioned into overlapping tiles of fixed size so memory and runtime remain bounded, which enables processing of dense streets without the erosion of thin structures that accompanies coarse voxel encodings. Change decisions are computed per object using 3D occupancy overlap, normal-direction displacement, height and volume differences, and a histogram distance in the intersected volume, and every statistic is gated by the local level-of-detection. This object-level formulation directly resolves Added and Removed when overlap is partial, avoids cell-wise artifacts from slope and curb geometry, and reduces false positives that arise from viewpoint or sampling change. The output is a per-instance change table and tiled maps with the five labels Added, Removed, Increased, Decreased, and Unchanged together with confidence, which are aggregated over the full extent to support HD-map maintenance and autonomous driving use cases.
Figure 2: Subiaco driving loops for the 2025 campaign. Colored polylines denote twenty overlapping loops that provide repeated viewpoints across arterials, residential blocks, and intersections.
Figure 3: Sensor rig and installation. The system comprises an Ouster OS1-128 LiDAR, a dual-antenna RTK GNSS with MEMS IMU, and a logging unit mounted on a reinforced roof plate with a measured antenna baseline.
@@
We validate the approach on bi-temporal city-scale maps of Subiaco by constructing a new 2025 high-definition LiDAR map and comparing it with the 2023 counterpart [37] under object-level annotations. Aggregated over 15 representative blocks, the method attains 95.2% ACC, 90.4% mF1, and 82.6% mIoU. Compared with the strongest baseline, Triplet KPConv, these results correspond to absolute gains of 0.2 percentage points in accuracy, 0.2 percentage points in mF1, and 0.8 percentage points in mIoU. The largest classwise improvement appears on Decreased, where the IoU reaches 74.8%, exceeding the SOTA baseline by 7.6 percentage points.
Our contributions are summarized below:
@@
    Dataset. A city-scale HD LiDAR dataset for Subiaco 2025 with a globally referenced .ply map, per-instance semantics for {ground, building, vegetation, mobile}, and object-level change labels obtained by registering the public 2023 map to the 2025 frame. Released on IEEE DataPort.
@@
    Method. An object-centric, uncertainty-aware 3D change detector that combines city-scale registration, local level-of-detection gating, class-consistent association, and instance-wise decisions on overlap, displacement, height, and volume. On 15 representative blocks it achieves 95.2% accuracy, 90.4% mF1, and 82.6% mIoU.
@@
    Pipeline. A reproducible end-to-end workflow for urban change mapping that spans acquisition, map construction, semantic and instance segmentation, association, and change analysis with bounded-memory tiling. Code and configurations are available at GitHub.
II Subiaco 2025 3D LiDAR Map Construction
This section specifies the acquisition and mapping workflow that yields the 2025 Subiaco reference used throughout. It covers survey design and sensor configuration, logging and georeferencing, packet deskewing and density standardization, intra and inter sequence alignment with loop closures and GNSS priors, and the final fusion and height normalization that produce a single globally referenced point cloud at uniform resolution. The goal is to make explicit the data fidelity and coordinate frame assumptions required by the downstream registration, segmentation, and change analysis stages.
@@ II Subiaco 2025 3D LiDAR Map Construction
II-A Survey Plan and Sensor Setup
The survey covered Subiaco with overlapping road loops that span arterial corridors, residential blocks, and major intersections to force repeated viewpoints for loop closure and cross-street alignment. Figure 2 illustrates the twenty loops used in 2025. Data were collected from a roof-mounted rig at 10 to 30 km/h during low-traffic windows. The primary sensor is an Ouster OS1-128 operated at 10 Hz in dual-return mode with a $45^{\circ}$ vertical field of view, $0.35^{\circ}$ angular resolution, and approximately 2.62 M points per second. Geodetic reference and motion state come from a dual-antenna RTK GNSS with a MEMS IMU. GNSS fixes are logged at 5 Hz and IMU at 73 Hz. Sensors are rigidly co-located on a reinforced roof plate with a measured antenna baseline to improve yaw observability, as shown in Figure 3. Hardware time is synchronized with GNSS PPS. The LiDAR accepts the PPS and stamps packets with synchronized timestamps. All streams are recorded on an Ubuntu host using GigE for LiDAR and USB or serial for GNSS and IMU. Factory extrinsics are refined by a short constrained-motion routine and verified by reprojection residuals on planar façades. Raw LiDAR packets (.pcap) are converted to per-scan .ply with dual returns and per-point timestamps, and GNSS or IMU logs are exported as .csv with UTC time, quality flags, and orientation.
II-B Acquisition Summary
Each route is logged as an independent sequence with continuous 10 Hz LiDAR frames, 5 Hz GNSS fixes (RTK state recorded), and 73 Hz IMU samples. LiDAR frames contain 260k–280k points per scan with intensity and ring index. We retain both Strongest and Last returns; mapping uses Strongest by default while both are archived for analysis. Sequences were recorded under dry daylight conditions to minimize rain and multipath artifacts; segments with RTK loss or packet drop exceeding pre-set thresholds are flagged in a YAML manifest. All trajectories and maps are georeferenced to GDA2020 / MGA Zone 50 and stored as globally referenced .ply maps together with per-sequence odometry in KITTI format.
II-C Map Generation
LiDAR packets are deskewed using per-pixel timestamps and the closest IMU segment. Statistical outlier removal and a voxel pyramid (1.0/0.5/0.25 m) standardize density. Within each sequence, multi-resolution NDT provides stable coarse alignment; point-to-plane ICP refines at each level. Loop closures are proposed by 360° LiDAR descriptors and verified by geometric consistency; a pose graph solves for globally consistent per-sequence poses. GNSS is used as a soft spatial prior: pose graph nodes that coincide with high-quality GNSS fixes (based on RTK status and dilution of precision) are anchored with adaptive weights, while GNSS-denied segments are down-weighted. Adjacent sequences are merged by aligning overlapping submaps with feature-initialized ICP and then fused. After global fusion, ground tiles are fit by robust plane models to define the vertical axis and zero height per tile; both epochs are rotated to align the vertical to $\mathbf{e}_{z}$ and are height-normalized to suppress long-wavelength bias. The final deliverable is a single globally referenced city-scale .ply map for 2025 at 0.25 m leaf size with per-point coordinates, intensity, and normals, which is the input to registration/LoD estimation and the semantic layer.
III Methodology
This section formalizes the object centric change detection pipeline on paired LiDAR maps. It introduces calibrated registration with a local level of detection to quantify uncertainty, derives geometry based correspondences, and constructs semantic and instance abstractions for ground, building, vegetation, and mobile classes. Class consistent association is then enforced and per object evidence from overlap, displacement, height, volume, and histogram cues is combined under uncertainty gating to produce five change labels with confidence. The design targets stability at city scale and supports reproducible evaluation and deployment. The end-to-end workflow and data flow are summarized in Fig. 4.
@@ III-A Geometric Registration and Object-Level Correspondence
Let $\mathcal{P}^{23}$ and $\mathcal{P}^{25}$ denote the city-scale LiDAR maps of Subiaco acquired in 2023 and 2025, respectively, each stored as a globally georeferenced .ply. This layer estimates a single rigid transform $T_{23\rightarrow 25}\!\in\!\mathrm{SE}(3)$ to align $\mathcal{P}^{23}$ onto $\mathcal{P}^{25}$, propagates registration uncertainty to a local level-of-detection (LoD), and derives object-level correspondences (LOLocation) that will be refined by the semantic layer.
We first apply statistical outlier removal and build a voxel pyramid with leaf sizes $\{1.0,0.50,0.25\}$ m to equalize density and reduce computation. Surface normals are estimated per level using a fixed-radius neighborhood ($r_{n}=0.8$ m). Coarse alignment $T_{0}$ is obtained, when external priors are insufficient, by feature-based registration on the $1.0$ m level: ISS keypoints, FPFH descriptors, and RANSAC with salient/support radii $\{1.5,2.0\}$ m and an inlier threshold of $0.6$ m (20k iterations).
Fine alignment proceeds from coarse to fine with Normal Distributions Transform (NDT) followed by point-to-plane ICP. NDT supplies a stable pose update at each level; ICP then reduces the residual geometric error using target normals. Given correspondences $\{(p_{i},q_{i},n_{i})\}$, we minimize
@@
where $w_{i}$ down-weights long-range pairs via a Huber kernel on Euclidean distance and $\rho_{\tau}(\cdot)$ is Tukey’s biweight with $\tau=0.3$ m. We solve (1) by Gauss–Newton with analytic Jacobians and reject ill-conditioned updates via a normal-direction Fisher-information test. To stabilize long, low-curvature structures, a plane-to-plane penalty is added at the finest level after RANSAC plane extraction on façades and ground:
$\mathcal{L}_{\text{pl2pl}}(R)=\lambda\sum_{k}\big\|R\,\bar{n}^{(23)}_{k}-\bar{n}^{(25)}_{k}\big\|_{2}^{2},\hskip 18.49988pt\lambda=0.1,$ (2)
@@
$\mathrm{LoD}_{95}=1.96\sqrt{\frac{\sigma_{n,23}^{2}}{N_{23}}+\frac{\sigma_{n,25}^{2}}{N_{25}}+\sigma_{\mathrm{reg}}^{2}},$ (3)
@@
To initialize object-level associations prior to semantics, we extract geometry-only proxies. Ground cells are removed, and connected components are computed on a $0.5$ m occupancy grid with morphological closing. For each component $o$ we compute a centroid $c(o)$, an oriented bounding box $\mathrm{OBB}(o)$ from PCA, the $95^{\text{th}}$ height quantile $h_{95}(o)$, and covariance eigen-features (linearity, planarity, sphericity). Cross-epoch association uses gated nearest neighbors with a shape-consistency cost
$\mathsf{cost}\!\left(o_{i}^{23},o_{j}^{25}\right)$ $=\alpha\,\lVert c_{i}-c_{j}\rVert_{2}^{2}$
$\qquad+\beta\,\Bigl(1-\mathrm{IoU}\!\left(\mathrm{OBB}_{i},\mathrm{OBB}_{j}\right)\Bigr)$
$\qquad+\gamma\,D_{\chi^{2}}\!\left(H_{i},H_{j}\right).$ (4)
where $H$ is a 10-bin height histogram. We set $(\alpha,\beta,\gamma)=(1,2,0.5)$ and accept a pair if $\|c_{i}-c_{j}\|_{2}<2$ m, $\mathrm{IoU}>0.1$, and $\mathsf{cost}<3.5$. Unmatched components are forwarded as Added/Removed candidates. Matched pairs form the LOLocation table $\pi:\mathcal{O}^{23}\!\rightarrow\!\mathcal{O}^{25}$ with attributes $(c,\mathrm{OBB},h_{95},\ldots)$. The semantic/instance segmentation module subsequently refines identities and resolves split/merge cases while preserving these geometric correspondences.

### v2

@@ Abstract
High-definition 3D city maps enable city planning and change detection which is essential for municipal compliance, map maintenance, and asset monitoring, including both built structures and urban greenery. Conventional Digital Surface Model (DSM) and image differencing are sensitive to vertical bias and viewpoint mismatch, while original point cloud or voxel models require large memory, assume perfect alignment, and degrade thin structures. We propose an uncertainty-aware, object-centric method for city-scale LiDAR-based change detection. Our method aligns data from different time periods using multi-resolution Normal Distributions Transform (NDT) and a point-to-plane Iterative Closest Point (ICP) method, normalizes elevation, and computes a per-point level of detection from registration covariance and surface roughness to calibrate change decisions. Geometry-based associations are refined by semantic and instance segmentation and optimized using class-constrained bipartite assignment with augmented dummies to handle split–merge cases. Tiled processing bounds memory and preserves narrow ground changes, while instance-level decisions integrate overlap, displacement, and volumetric differences under local detection gating. We perform experiments on the city of Subiaco (Western Australia) dataset, captured once in 2023 and again in 2025. Our method achieves 95.3 % accuracy, 90.8 % mF1, and 82.9 % mIoU, improving over the strongest baseline, Triplet KPConv, by 0.3, 0.6, and 1.1 points, respectively. The datasets are available on IEEE DataPort (2023 and 2025). The source code has been released on Github Repository.
@@ I Introduction
Modern cities are implementing smart transportation, digital twins, and autonomous driving systems that rely on high-definition (HD) 3D city maps maintained at an operational mapping cadence [1, 2]. This creates the need for change detection i.e. identifying additions, removals, and volumetric variations across bitemporal 3D HD maps. Besides HD map maintenance, change detection is also useful for construction monitoring, municipal compliance, asset management, and timely updates for localization and planning in autonomy stacks [3, 4, 5, 6, 7]. LiDAR (Light Detection And Ranging) sensor supports this process by providing dense 3D geometry–by driving a LiDAR mounted car through the city–that distinguishes long-term structural changes from transient dynamics such as vehicles and pedestrians [8, 9, 10].
@@
Figure 1: Overview of the proposed 3D change-detection pipeline. Analysing cross-epoch association, semantic and instance segmentation, and uncertainty-gated labels.
Existing approaches to 3D change detection can be grouped into height- or DSM-based differencing, 2D image-based methods, and point-wise deep learning models [11, 12]. Height differencing demands precise co-registration under uniform sampling [13, 14]. In dense streets, small vertical biases and ground slopes cause systematic errors, while curbs and facade parallax yield false positives; moreover, results remain cell-wise without object identity [15, 13]. Image-based methods cannot resolve geometry and are vulnerable to viewpoint inconsistencies across satellites and time periods [16, 17]. Point-wise Siamese and triplet networks, such as KPConv variants, process cropped tiles with millions of points [18, 19, 20]. Their runtime and memory scale with area, training assumes balanced labels, and per-point predictions weaken Added/Removed reasoning when overlap is partial or occlusion alters sampling [21, 22]. Voxel or bird’s-eye view encodings reduce memory but erode thin structures and fine-scale ground changes, forcing a trade-off between coverage and fidelity [23, 24].
Many pipelines also assume perfect pre-alignment and fail to propagate registration uncertainty into decision thresholds, leading to spurious detections near loop closures or under a canopy [25, 26, 27, 28, 29]. Furthermore, few methods enforce class-consistent associations between epochs, leaving split and merge cases unresolved and causing per-class count drift at city scale [30, 31, 32]. These constraints limit deployment for operational HD-map maintenance in dense urban corridors and highlight the need for a city-scale, object-centric, and uncertainty-aware approach.
We propose a change detection method that addresses the above limitations. Our method is object-centric and uncertainty-aware, operating at city scale on bi-temporal LiDAR maps.
Figure 2: Subiaco driving loops for the 2025 data capture. Colored polylines denote twenty overlapping loops for individual reconstruction and merging. When comparing change detection results with other learning based methods, we take the mutually exclusive parts of Loop 1 to 5 for testing and the rest for training.
An overview of our method is shown in Fig. 1. The process begins with global alignment using multi-resolution NDT [33] followed by point-to-plane ICP [34]. Both maps are then height-normalized, and a per-location level of detection is derived from registration covariance and surface roughness. This replaces the common assumption of perfect pre-alignment and introduces calibrated thresholds that suppress spurious detections along loops and under canopy. Geometry-only proxies initiate cross-epoch associations, followed by semantic and instance segmentation for ground, building, vegetation, and mobile classes on single-epoch geometry. Class-constrained bipartite matching refines correspondences and resolves split and merge cases through augmented assignment with dummies, thereby preventing per-class count drift and enforcing label consistency over time.
The city is partitioned into overlapping tiles of fixed size to bound memory and runtime, enabling dense street processing without erosion of thin structures that typically occurs with coarse voxel encodings. Change decisions are made per object using 3D occupancy overlap, normal displacement, and volumetric differences, each gated by the local level of detection. This object-level formulation resolves Added and Removed cases under partial overlap, mitigates slope- or curb-induced artifacts, and reduces false positives caused by viewpoint or sampling variations. The output is a per-instance change table and tiled maps labeled Added, Removed, Increased, Decreased, or Unchanged, each with confidence, aggregated city-wide to support HD-map maintenance and autonomous driving applications.
Another bottleneck in advancing outdoor city-scale change detection research is the lack of public benchmarks. Existing public benchmarks for 3D change detection cover limited area or rely on aerial data, and object-level annotations across data captured at different times remain scarce [35, 32]. Some of the datasets and methods work on simulated point clouds rather than detecting actual changes that occur over long periods of time [32]. We also address this gap by releasing a follow-on Subiaco (Western Australia) city-scale dataset that extends our 2023 capture [36]. The new dataset was acquired in 2025, after a two-year interval. Subiaco, immediately west of Perth’s central business district, serves as a major transport hub with two arterial roads and heavy daily commuting, making it a representative urban testbed for long-term change analysis [37].
@@
We validate the proposed method on the bi-temporal city-scale maps of Subiaco from 2023 and 2025. A new 2025 HD LiDAR map is constructed and compared with the public 2023 counterpart [36] under object-level annotations. Across 15 representative blocks, the method attains 95.3% accuracy, 90.8% macro F1, and 82.9% macro IoU. When compared to the strongest baseline, Triplet KPConv, the improvements are 0.3 points in accuracy, 0.6 points in macro F1, and 1.1 points in macro IoU. The largest class-wise gain occurs for Decreased, where the IoU reaches 75.6%, exceeding the baseline by 8.4 points.
Our main contributions are as follows:
@@
    Method: We propose an object-centric, uncertainty-aware 3D change detection method that integrates city-scale registration, local detection gating, class-consistent association, and instance-wise metrics on overlap, displacement, height, and volume. On 15 representative blocks from the Subiaco dataset, it achieves 95.3% accuracy, 90.8% mean F1, and 82.9% mean IoU.
@@
    Dataset: We propose a city-scale HD LiDAR dataset for city-scale change detection. Our new Subiaco 2025 dataset includes a globally referenced .ply map, per-instance semantics for {ground, building, vegetation, mobile}, and object-level change labels obtained by registering our new 2025 map to the public 2023 Subiaco map [36]. The 2025 Subiaco City Map datasets are available on IEEE Dataport.
@@
    End-to-end Workflow: We propose a reproducible end-to-end workflow for urban change mapping encompassing acquisition, map construction, semantic and instance segmentation, association, and change analysis with bounded-memory tiling. The source code has been released on Github Repository.
Figure 3: Multi-sensor rig. The system comprises an Ouster OS1-128 LiDAR, a dual-antenna RTK GNSS with MEMS IMU, and a logging unit mounted on a reinforced roof plate with a measured antenna baseline.
II Data Capture and 3D City Map Construction
This section specifies the data acquisition and mapping workflow to generate the 3D map of the city of Subiaco in Western Australia. We refer to it as the 2025 Subiaco map for short. This section covers survey design (see Fig 2) and sensor configuration, logging and georeferencing, packet deskewing and density standardization, intra and inter sequence alignment with loop closures and GNSS priors, and the final fusion and height normalization to produce a single globally referenced point cloud at uniform resolution. The goal is to make explicit the data fidelity and coordinate frame assumptions required by the downstream registration, segmentation, and change analysis stages.
II-A Survey Plan and Multi-Sensor Setup
The survey covers Subiaco with overlapping road loops that span arterial corridors, residential blocks, and major intersections to force repeated viewpoints for loop closure and cross-street alignment. Fig 2 illustrates the twenty loops used in 2025. Data were collected from a roof-mounted rig on a car driving at 10 to 30 km/h during low-traffic windows. The primary sensor is an Ouster OS1-128 operated at 10 Hz in dual-return mode with a $45^{\circ}$ vertical field of view, $0.35^{\circ}$ angular resolution, recording approximately 2.62 M points per second. Geodetic reference and motion state come from a dual-antenna RTK GNSS with a MEMS IMU. GNSS fixes are logged at 5 Hz and IMU at 73 Hz. Sensors are rigidly mounted on a reinforced roof top rack with a measured antenna baseline to improve yaw observability, as shown in Fig 3. Hardware time is synchronized with GNSS PPS. The LiDAR accepts the PPS and stamps packets with synchronized timestamps. All streams are recorded on an Ubuntu host using GigE for LiDAR and USB cable for GNSS and IMU. Factory extrinsics are refined by a short constrained-motion routine and verified by reprojection residuals on planar facades. Raw LiDAR .pcap packets are processed into per-scan .ply files containing dual returns and per-point timestamps. GNSS and IMU logs are exported to .csv format with UTC timestamps, quality indicators, and orientation measurements.
II-B Data Acquisition Summary
Each route is logged as an independent sequence with continuous 10 Hz LiDAR frames, 5 Hz GNSS fixes (RTK state recorded), and 73 Hz IMU samples. LiDAR frames contain 260k–280k points per scan with intensity and ring index. We retain both strongest and last returns; mapping uses strongest by default while both are archived for analysis. Sequences were recorded under dry daylight conditions to minimize rain and multipath artifacts; segments with RTK loss or packet drop exceeding pre-set thresholds are flagged in a YAML manifest. All trajectories and maps are georeferenced to GDA2020 / MGA Zone 50 and stored as globally referenced .ply maps together with per-sequence odometry, similar to the KITTI format [38].
Figure 4: Pipeline of acquisition and mapping (soft GNSS synchronization, ground segmentation, sequence construction, NDT followed by ICP). Right:
II-C 3D City Map Generation
Fig 4 shows an overview of the map generation. LiDAR packets are deskewed using per-pixel timestamps and the closest IMU segment. Statistical outlier removal and a voxel pyramid (1.0/0.5/0.25 m) standardize the density. Within each sequence, multi-resolution NDT [33] provides stable coarse alignment; point-to-plane ICP [34] at each level. Loop closures are proposed by 360° LiDAR descriptors and verified by geometric consistency; a pose graph solves for globally consistent per-sequence poses. GNSS is used as a soft spatial prior: pose graph nodes that coincide with high-quality GNSS fixes (based on RTK status and dilution of precision) are anchored with adaptive weights, while GNSS-denied segments are down-weighted. Adjacent sequences are merged by aligning overlapping submaps with feature-initialized ICP and then fused. After global fusion, ground tiles are fit by robust plane models to define the vertical axis and zero height per tile; both epochs are rotated to align the vertical to $\mathbf{e}_{z}$ and are height-normalized to suppress long-wavelength bias. The final deliverable is a single globally referenced city-scale .ply map for 2025 at 0.25 m leaf size with per-point coordinates, intensity, and normals, which is the input to registration/level of detection (LoD) estimation and the semantic layer.
@@ II-B Data Acquisition Summary
III Change Detection Method
This section formalizes the proposed object centric change detection method on paired LiDAR maps. It introduces calibrated registration with a local level of detection to quantify uncertainty, derives geometry based correspondences, and constructs semantic and instance abstractions for ground, building, vegetation, and mobile classes. Class consistent association is then enforced and per object evidence from overlap, displacement, height, volume, and histogram cues is combined under uncertainty gating to produce five change labels with confidence. The design targets stability at city scale and supports reproducible evaluation and deployment. The proposed end-to-end workflow and data processing are summarized in Fig. 5.
@@ III-A Geometric Registration and Object-Level Correspondence
Let $\mathcal{P}_{2023}$ and $\mathcal{P}_{2025}$ denote the city-scale LiDAR maps of Subiaco acquired in 2023 and 2025, respectively, each stored as a globally georeferenced .ply. This layer estimates a single rigid transform $T_{2023\rightarrow 2025}\!\in\!\mathrm{SE}(3)$ to align $\mathcal{P}_{2023}$ onto $\mathcal{P}_{2025}$, propagates registration uncertainty to a local level-of-detection, and derives level object location that will be refined by the semantic layer.
We first apply statistical outlier removal and build a voxel pyramid with leaf sizes $\{1.0,~0.50,~0.25\}$ m to equalize density and reduce computation. Surface normals are estimated per level using a fixed-radius neighborhood ($r_{n}=0.8$ m). Coarse alignment $T_{0}$ is obtained, when external priors are insufficient, by feature-based registration on the $1.0$ m level: ISS keypoints, FPFH descriptors, and RANSAC with salient/support radii $\{1.5,2.0\}$ m and an inlier threshold of $0.6$ m (20k iterations).
Precise alignment proceeds from coarse to fine with Normal Distributions Transform (NDT) followed by point-to-plane ICP. NDT supplies a stable pose update at each level; ICP then reduces the residual geometric error using target normals. Given correspondences $\{(p_{i},q_{i},n_{i})\}$, we minimize
@@
where $w_{i}$ down-weights long-range pairs via a Huber kernel on Euclidean distance and $\rho_{\tau}(\cdot)$ is Tukey’s biweight with $\tau=0.3$ m. We solve (1) by Gauss–Newton with analytic Jacobians and reject ill-conditioned updates via a normal-direction Fisher-information test. To stabilize long, low-curvature structures, a plane-to-plane penalty is added at the finest level after RANSAC plane extraction on facades and ground:
$\mathcal{L}_{\text{pl2pl}}(R)=\lambda\sum_{k}\big\|R\,\bar{n}^{(2023)}_{k}-\bar{n}^{(2025)}_{k}\big\|_{2}^{2},\qquad\lambda=0.1,$ (2)
@@
$\mathrm{LoD}_{95}=1.96\sqrt{\frac{\sigma_{n,2023}^{2}}{N_{2023}}+\frac{\sigma_{n,2025}^{2}}{N_{2025}}+\sigma_{\mathrm{reg}}^{2}},$ (3)
@@
To initialize object-level associations prior to semantics, we extract geometry-only proxies. Ground cells are removed, and connected components are computed on a $0.5$ m occupancy grid with morphological closing. For each component $o$ we compute a centroid $c(o)$, an Oriented Bounding Box (OBB) $\mathrm{OBB}(o)$ from Principal Component Analysis (PCA), the $95^{\text{th}}$ height quantile $h_{95}(o)$, and covariance eigen-features (linearity, planarity, sphericity). Cross-epoch association uses gated nearest neighbors with a shape-consistency cost
$\begin{split}s_{y}(o_{i},o_{j})&=\Bigl(1+\exp\!\bigl[-\bigl(w_{h}z_{h}+w_{v}z_{v}+w_{o}z_{o}+w_{c}z_{c}\\ &\qquad\qquad\qquad{}+\,w_{\perp}\,\tfrac{|D_{\perp}(o_{i},o_{j})|}{\overline{\mathrm{LoD}}_{95}}\bigr)\bigr]\Bigr)^{-1}.\end{split}$ (4)
Here $c$ is the object centroid, $\mathrm{IoU}$ the 3D box overlap ratio, $H$ a 10-bin height histogram per object, and $D_{\chi^{2}}$ the chi-square distance between histograms. We set $(\alpha,\beta,\gamma)=(1,2,0.5)$ and accept a pair if $\|c_{i}-c_{j}\|_{2}<2$ m, $\mathrm{IoU}>0.1$, and $\mathsf{cost}<3.5$. Unmatched components are treated as Added/Removed. Accepted pairs define $\pi:\mathcal{O}_{2023}\!\rightarrow\!\mathcal{O}_{2025}$ with attributes $(c,\mathrm{OBB},h_{95})$.